what is the enthalpy of a gas at 20 K given that its heat capacity as a function of temperature is given by C=2T^2, over the range T= 0K to 20K ?


1
Expert's answer
2021-05-18T13:53:44-0400

At constant pressure, the heat capacity is


"C_P=\dfrac{dH}{dT}"


Then

"dH=C_PdT"

Integrate both sides


"\Delta H=\displaystyle\int_{T_i}^{T_f}C_PdT"



Given "C_P=2T^2, T_i=0K, T_f=20K"



"\Delta H=\displaystyle\int_{0}^{20}2T^2dT=\bigg[\dfrac{2T^3}{3}\bigg]\begin{matrix}
 20 \\
 0
\end{matrix}"

"=\dfrac{16}{3}\times10^3 \ J"